OpenStudy (anonymous):

If x - 3 = 12, then what is the value of x - 3 + 3? Why?

OpenStudy (anonymous):

x-3=12 take the -3 across to become +3 x=12+3 x=15 the -3 and +3 cancel so the value is 15